Please check my work, am I doing this correctly?
Solve using the square root property:
x+8)^2=81
Þ ((x+8)2-81) = 0
Þ (x2+16x-17) = 0
Þ (x2-x+17x-17) = 0
Þ x(x-1)+17(x-1) = 0
Þ (x+17)(x-1) = 0
Þ (x+17) = 0 OR (x-1) = 0
Þ x = -17 OR x = 1

Answer by priya(51)

There is another simple step to get ur answer.
the given equation is (x+8)^2 = 81

this can be written as (x +8)^2 = 9^2 (9*9= 81 right)
now taking sqrt on both sides we get,

x +8 =+ OR - 9
therefore FOR +9, x = 9 -8
X =1
FOR -9, x= -9 -8
x = -17
IS THIS SIMPLE..
YEAH..
VISIT MY SITE (CLICK ME) TO LEARN MATHS ONLINE..
STEP BY STEP TEACHING..
PRIYA